OpenStudy (anonymous):

Help!! So confused! Dante is making a necklace with 18 rows of tiny beads in which the number of beads per row is given by the series 3 + 10 + 17 + 24 + ... a. Use summation notation to write the series. Explain what the numbers in the summation notation represent in this situation and how you found the expression used in the summation. b. Find the total number of beads in the necklace. Explain your method for finding the total number of beads.

OpenStudy (anonymous):

find the difference between the successive terms...

OpenStudy (anonymous):

notice a trend there?

OpenStudy (anonymous):

They each go up by 7

OpenStudy (anonymous):

what is such a series called?

OpenStudy (anonymous):

arithmetic

OpenStudy (anonymous):

good job.. so you got the answer for (a) now?

OpenStudy (anonymous):

How would I write the summation notation for question a?

OpenStudy (anonymous):

ok, start by writing what we have.. 3 = 3 + 0*7 10 = 3 + 1*7 17 = 3 + 2*7 and so on 18 times a_n = a_0 + (n-1)*d that give you the n'th number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now, all you gotta do is add em up \[ S = \sum_{n=1}^{18}a_n\\ S=\sum_{n=1}^{18}a_0+(n-1)d \]

OpenStudy (anonymous):

where a0=? d=?

OpenStudy (anonymous):

d= the common difference so d would = 7

OpenStudy (anonymous):

and a would be the 1st term so would that = 3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yay! Okay so to find the total number of beads which is question #b I would multiply?

OpenStudy (anonymous):

no.. total mean a sum... 3 apples in one hand and 4 apples in the other.. total is? (add)

OpenStudy (anonymous):

so, the sum of arithmatic series to nth term is: \[S_n={n\over2}[2a_0+(n-1)d]\]

OpenStudy (anonymous):

how many rows you have? 18.. so, n = 18 and rest you already have

OpenStudy (anonymous):

ohh ok. so would the answer be 1125 beads?

OpenStudy (anonymous):

i gues.. (9)(6+17*7)